OpenStudy (anonymous):

2x+5. when x=10? help.

OpenStudy (anonymous):

2(10)+5 20+5 = 25

OpenStudy (anonymous):

replace x with 10 and solve.

OpenStudy (k_lynn):

Ok, @lolowlet since x=10, plug that number into your equation where the x's are. So now your equation will look like this >2*10+5=n. Let n represent your answer. We're going to work according to PEMDAS. There are no parentheses, or exponents, but we do have some multiplication in there, solve 2*10=n for the answer of 20. So n=20. Remember that number because you'll need it later. There is no division in the equation but there is some addition, so now we're going to add 20+5=y. Let y represent the answer. You should get 25. That is the answer. Hope this helped